# Technical insights and legal considerations for advancing federated learning in bioinformatics

### v1

@@ 5 Legal aspects of federated learning
The legal frameworks used within FL consortia are rarely discussed in the literature. Ballhausen et al. (2024) describes both the technical and legal aspects of a European pilot study implementing a federated statistical analysis by secure multiparty computation. They set up agreements between parties similar to those between the participants of a multi-centre clinical trial because using SMPC and only exchanging model gradients was legally considered data pseudonymisation (rather than anonymisation). FL was determined to require the same level of data protection as regular data sharing, which is also the most conservative course of action suggested in Truong et al. (2021); Lieftink et al. (2024). All clients jointly controlled the consortium and were responsible for determining the purpose and means of processing, including obtaining approval from the respective Ethics Committees. Sun et al. (2022a) similarly describes the server in their consortium as a trusted, secure environment supported by a legal joint controller agreement between the data owners.
@@
Following Ballhausen et al. (2024), establishing an FL consortium could be expected to require all participating and involved parties to execute agreements that regulate their interactions. Doing so will establish the level of trust between parties and their responsibilities towards each other, third parties, and patients. Risk aversion suggests that it should contain a data-sharing agreement allowing for information sharing described in Figure 1. No party has access to other parties’ data. Still, it is theoretically possible that, in some cases, the model updates shared during FL could be deanonymised by malicious internal or external attackers (Truong et al., 2021). Parties may then be reluctant to treat that information as non-personally identifiable without formal mathematical proof of anonymisation and prefer to establish data protection responsibilities with a data-sharing agreement. In the EU, GDPR states that we should consider “all the means reasonably likely to be used” to “determine whether a natural person is identifiable” (European Union, 2016). Securing infrastructure in depth using best practices from information technology, defensive software engineering, and data by secure computing and encryption can make malicious attacks impractical with current technologies (security by design and by default; Volini, 2020; Martin and Kung, 2018). In addition, when FL involves models other than deep neural networks, if the contributions of individual parties are well balanced across the consortium and include a sufficiently large number of individuals, the information exchanged may very well be the same summary information routinely published as supplementary material to academic journal publications (Jégou et al., 2024). A recent systematic literature review of privacy attacks in FL has also highlighted that many of them are only feasible under unrealistic assumptions (Wainakh et al., 2022). Therefore, the reduced amount of information shared during FL and using secure computing should be considered to provide increased protection against data leaks and misuse. Advertising it as a key feature of the FL consortium will make partners and patients more comfortable with contributing to federated studies (see, for instance, Ballhausen et al., 2024). Lieftink et al. (2024), which investigated how FL aligns with GDPR in public health, also acknowledges that FL mitigates many privacy risks by enforcing purpose limitation, data use and information exchange minimisation, integrity and confidentiality (at a cost, as discussed in Section 2.4).
@@
Furthermore, consortium parties must agree on how to assign intellectual property (IP) rights. Bioinformatics research often has practical applications in industry, which may involve patenting the results and apportioning any financial gains arising from their use. Parties in the consortium jointly control it and should share any gains from it (Minssen and Pierce, 2018). FL consortia are no different in this respect. From a technical standpoint, watermarking techniques for tracking data provenance and plagiarism have been adapted to FL Tekgul et al. (2021) to identify data and model theft.
@@
In addition, the agreement establishing the FL consortium must set out its relationships with third parties and their legal duties. Third parties that have access to the infrastructure may be required to sign a data processing agreement to guarantee the safety and privacy of data. In many countries, patients also have the right to withdraw their consent to use their data, which in turn may require implementing procedures to remove individual data points from future federated analyses.
@@
To summarise, FL provides increased protection against data and model leaks and should reduce the parties’ and patients’ perceived risk in contributing to the consortium. However, out of an abundance of caution, establishing a consortium-wide data-sharing agreement may allocate and reduce party responsibilities in the event of privacy breaches. The use of FL has a limited impact on other legal aspects of collaborative analysis, such as IP handling and requirements for third parties, because it is a technical solution that does not change the fundamental legal rights and responsibilities of the parties involved in the consortium.
@@ 6 Conclusions
Several independent research efforts have shown federated learning to be an effective tool to improve clinical discovery while minimising data sharing (Roth et al., 2020; Pati et al., 2022; Heyndrickx et al., 2023). FL allows access to larger and more varied data pools, leading to a faster and more robust exploration and translation of results. At the same time, it provides enhanced data privacy and can seamlessly incorporate advanced encrypted and secure computation techniques. Despite the increased computational requirements and reduced ability to explore and troubleshoot issues with the data (Lieftink et al., 2024), the benefits of FL may outweigh these additional costs.
Therefore, federated learning can potentially reduce the risks associated with EU and US regulations if implemented in a manner secure by design and by default. Its use may also make patients and institutions more confident in participating in clinical studies by reducing privacy and data misuse risks. However, the reliable use of federated learning and its effective translation into clinical practice requires a concerted effort by machine learning and clinical and information technology specialists. All their skills are required to accurately evaluate those risks and expand practical applications of federated learning in bioinformatics beyond the early-stage applications reviewed in this paper.

### v4

@@ 6 Legal aspects of federated learning
The legal frameworks used within FL consortia are rarely discussed in the literature. Ballhausen et al. (103) describes both the technical and legal aspects of a European pilot study implementing a federated statistical analysis by secure multiparty computation. They established agreements between parties similar to those between participants in a multi-centre clinical trial, as using SMPC and exchanging model gradients was legally considered data pseudonymisation (rather than anonymisation). FL was determined to require the same level of data protection as regular data sharing, which is also the most conservative course of action suggested in Truong et al. (9), Lieftink et al. (104). All clients jointly controlled the consortium and were responsible for determining the purpose and means of processing, including obtaining approval from the respective Ethics Committees. Sun et al. (105) similarly describes the server in their consortium as a trusted and secure environment, supported by a legal joint controller agreement between the data owners.
@@
Following Ballhausen et al. (103), establishing an FL consortium could be expected to require all participating and involved parties to execute agreements that regulate their interactions, the so-called DPA or DSA (data protection/processing/sharing agreements). Doing so will establish the level of trust between parties and their responsibilities towards each other, third parties, and patients. Risk aversion suggests that it should include a data-sharing clause to allow for the sharing of information, similar to a centralised analysis, as described in Figure 1. No party has access to the data of other parties. Still, it is theoretically possible that, in some cases, the model updates shared during FL could be deanonymised by malicious internal or external attackers (9). Parties may then be reluctant to treat that information as non-personally identifiable without formal mathematical proof of anonymisation and prefer to establish data protection responsibilities with a data-sharing agreement. In the EU (GDPR), but also other jurisdictions (national data protection laws), “all the means reasonably likely to be used should be considered to determine whether a natural person is identifiable” (1). Securing infrastructure in depth using best practices from information technology, defensive software engineering, and data by secure computing and encryption can make malicious attacks impractical with current technologies (security by design and by default; 106, 107). In addition, when FL involves models other than deep neural networks, if the contributions of individual parties are well balanced across the consortium and include a sufficiently large number of individuals, the information exchanged may very well be the same summary information routinely published as supplementary material to academic journal publications (108). A recent systematic literature review of privacy attacks in FL has also highlighted that many of them are only feasible under unrealistic assumptions (8). Therefore, reducing the amount of information shared during FL and using secure computing must be considered to provide increased protection against data leaks and misuse. Advertising such measures as a key feature of the FL consortium will make partners and patients more comfortable with contributing to federated studies (see, for instance, 103). Lieftink et al. (104), which investigated how FL aligns with GDPR in public health, also acknowledges that FL mitigates many privacy risks by enforcing purpose limitation, data use and information exchange minimisation, integrity and confidentiality (at a cost, as discussed in Section 3.4).
@@
Furthermore, consortium parties must agree on how to assign intellectual property (IP) rights. Bioinformatics research often has practical applications in industry, which may involve patenting the results and apportioning any financial gains arising from their use. Parties in the consortium jointly control it and should share any gains from it (109). FL consortia are no different in this respect. From a technical standpoint, watermarking techniques for tracking data provenance and plagiarism have been adapted to FL (110) to identify data and model theft.
@@
Additionally, the agreement establishing the FL consortium must outline its relationships with third parties and their corresponding legal obligations. Third parties that have access to the infrastructure may be required to sign a data processing agreement to guarantee the safety and privacy of data. In many countries in Europe, as well as in the US, patients have the right to withdraw their consent to use their data at any time. This, in turn, may require implementing procedures to remove individual data points from future federated analyses.
@@
We summarised these considerations in Table 1, along with the key differences from the alternatives we discussed in Section 3. FL provides increased protection against data and model leaks, which should reduce the perceived risk for parties and patients in contributing to the consortium. However, out of an abundance of caution, establishing a consortium-wide data-sharing agreement may help allocate and reduce party responsibilities in the event of a privacy breach. The use of FL has a limited impact on other legal aspects of collaborative analysis, such as IP handling and requirements for third parties, because it is a technical solution that does not change the fundamental legal rights and responsibilities of the parties involved in the consortium. Table 3 summarises the key legal and procedural steps required to implement federated learning in biomedical research, as detailed above.
| Action | Documentation | Data controllers responsabilities | Data processors responsibilities | Notes |
| Ethical approval | Study protocol (including data sharing information). | Obtain approval from local ethics committees. | – | Making the use of secure computing transparent (thus limiting the ways data can be processed) should support ethics committees’ trust in the approach. |
| Consortium setup | Cooperation/ data protection/ processing/ sharing agreements | Joint controllers, responsible for the purposes and means of processing and for data security and purpose adherence. | Contractual responsibilities stemming from controllers’ data security and purpose adherence responsibilities. | Although parties cannot access each other’s data, agreements often permit the sharing of private information to address potential leakages from shared model parameters. |
| Data Protection Officer (DPO) advice | GDPR-compliant documentation, Data Protection Impact Assessment (DPIA), software documentation. | Appoint DPO if needed under GDPR (Article 35) and obtain advice. | - | Consortium agreements should specify how DPO responsibilities are allocated; a single DPO may be designated at the consortium level or appointed from a subset of partners. |
| Clinical data collection (retrospective or prospective) | Informed consent forms, data collection protocols. | Collect consent from all patients. Guarantee the right of withdrawal. | – | Data should generally be considered pseudo-anonymised. Explaining secure computing methods to patients should foster trust and informed participation. |
| Code deployment | Software license, deployment agreements. | Grant local software deployment compliant with deployment agreements. | Guarantee the software and platform’s security and usage comply with the purposes and licences. | The consortium should agree on where deployment happens (e.g., trusted execution environment - TEE), what is deployed and how deployments are authorised. |
| Intellectual Property (IP) | IP ownership agreements, licensing terms, publication policies. | If the trained model is protected by IP rights, its ownership and usage are governed by the terms of the contract. | Do not automatically participate in IP generated from the trained model. | In FL, the trained global model is typically viewed as a joint IP artefact. There is ongoing work to establish IP allocation models, licensing templates, and tailored governance mechanisms. |
| Model Governance | Model versioning logs, audit trails, validation reports, access control policies. | Guarantee model trustworthiness | Facilitate model governance and auditability | While stringent requirements are requested only in case of model deployment, analytical or research-only contexts still require principled governance to ensure reproducibility, accountability, and ethical compliance. |
We now proceed to discuss how FL facilitates compliance with the key requirements of GDPR (1) through its architecture. Data providers, which have a complete control over and a more intimate knowledge of the data they collected as well as a direct connection to data subjects, act in a ”data controller” role (Articles 4 and 24), taking ”appropriate technical and organisational measures” (Article 25) to ensure privacy and security, thus minimising the risk of data breaches. Therefore, they can directly scrutinise their use, notify data subjects about it to request consent (Article 9); allow them to withdraw their data (Article 7); ensure lawful, fair and transparent processing (Articles 12–15); and directly assess risks to data subjects and minimise them through appropriate legal agreements. Consortium parties, which include both data controllers (as clients) and data processors (as servers, compute facilities, as defined in Article 4), are also required to use the techniques described in Section 3.5 to ensure data security and privacy beyond what is provided by base FL (privacy by design, Articles 24, 25 and 32). For the same reasons, FL facilitates compliance with the EU AI Act (2). Some of its requirements strengthen those in the GDPR, such as data minimisation, localisation, transparency, auditability, security, and data quality. Additionally, the EU AI Act requires efforts to mitigate bias, ensure the robustness of models, implement human oversight, and assess high-risk systems. Data controllers are in the best position to ensure these requirements are met. Collectively, they can provide more representative samples that are less prone to bias and fairness issues. Finally, the presence of multiple data controllers in the consortium also implies that these requirements are verified by several independent parties.
In contrast, the US AI Act is more flexible in its requirements, which are left to sector-specific agencies to define and enforce. It focuses more on party self-regulation and harm remediation rather than universal legal mandates and prevention. As a result, cross-border EU-US consortia should rely on the EU-US Data Protection Framework (111) or put in additional safeguards as required by the GDPR (Article 46). Even so, FL’s privacy stance naturally fits well with the practical implementation of the US AI Act.
7 Challenges and future directions
Federated learning (FL) has evolved rapidly over its relatively short lifetime, becoming a widely adopted methodology across diverse domains. In a recent article (112), which includes several authors of the seminal FL paper (6), the progress of the field is reviewed, and key challenges for its future development are outlined. The authors propose a refined definition of FL centred on privacy principles and analyse how core concepts such as data minimisation, anonymisation, transparency and control, verifiability, and auditability have evolved and are expected to play a major role in the future. They identify three primary challenges for the field: scaling FL to support large and multimodal models, overcoming operational difficulties arising from device heterogeneity and synchronisation constraints, and addressing the current lack of verifiability in deployed systems. As a potential means to address this latter aspect, the authors highlight trusted execution environments (TEEs) (113) as a promising technology. A TEE is a secure area within a processor that executes code and processes data in isolation from other software, protecting it from tampering or unauthorised access even if the main operating system is compromised.
Complementing these insights, other surveys examine additional aspects of the anticipated future developments in federated learning (FL) research. Wen et al. (114) call for more efficient encryption schemes and greater overall efficiency in FL, including strategies to reduce communication costs between clients and servers. They also emphasise the importance of novel aggregation strategies that can better handle client heterogeneity. In the context of multimodal FL, aggregation must often reconcile contributions when clients supply only partially overlapping modalities, which represents an additional complexity beyond standard heterogeneity. Yurdem et al. (115) highlight the emerging paradigm of FL as a Service, in which federated learning training is conducted through ready-to-use platforms such as FeatureCloud (86) and sfkit (19), discussed in Section 4.6. This approach allows institutions to participate in collaborative model development with minimal software and deployment effort, thereby simplifying implementation and facilitating cross-organisational collaboration. Looking ahead, the emergence of federated foundation models is expected to define the next phase of research in this field. Their development will require progress across several key dimensions, including improving efficiency through advanced aggregation methods and optimised computational and communication frameworks; strengthening trustworthiness by increasing robustness to attacks; and enhancing incentive mechanisms that reward clients according to the quality of the models provided. Collectively, these advances are expected to be essential for making large-scale federated models practical and scalable while maintaining manageable communication costs (116).
The prospects of federated learning in bioinformatics depend on how the legal and technical landscape will develop. Further legislation will progressively regulate the use of machine learning and AI models, defining and restricting how data can be shared and used. As its effects percolate through protocol and product development, many aspects of federated learning will likely take a more definite shape.
Firstly, the security and privacy risks are likely to become more clearly defined. Jurisprudence will naturally shift from general guiding principles, such as the EU AI Act, to practical compliance rule sets as products based on federated learning enter the market. How to assess sensitive aspects of data (re)use will also likely be standardised (117), using the vast collection of available data and model cards as a starting point (118, 119). What threat models are relevant, what security measures are appropriate at the infrastructure level, and what attacks are feasible will then become clear, possibly confirming the irrelevance of many that have been speculated in the literature (8). The evolution of encryption and differential privacy techniques may also allow different types of data to be treated as anonymised, depending on their nature and the theoretical guarantees of those techniques. Some data (say, single-cell transcriptomics) are intrinsically more difficult to tie to a specific individual than others (say, whole genome sequences), and different privacy-enhancing techniques are more effective than others at mitigating various types of risk.
Secondly, the evolution of data and models will require periodic reevaluation of the trade-offs discussed in this paper. The ever-increasing volumes of data used to train federated models will eventually force a cost-benefit analysis for resource-intensive techniques like HE and MPC; lighter approaches may be the only feasible choice at scale, provided regulations permit their use. Modelling approaches will undergo a similar selection process, as they will be required to evolve to handle new biological and clinical data types in larger quantities and across multiple modalities. In this respect, we foresee that federated learning in bioinformatics may diverge from the mainstream, which has largely standardised on deep learning models (15, 16). Historically, bioinformatics has produced distinct model classes for different types of data. Federating them, optimising them, and assessing their privacy risks will require significant research and engineering efforts before they are suitable for practical applications.
@@ 8 Conclusions
Independent research efforts in several bioinformatics domains have shown federated learning to be an effective tool to improve clinical discovery while minimising data sharing (10, 12, 13, among others). FL enables access to larger and more diverse data pools, resulting in faster and more robust exploration and interpretation of results. At the same time, it provides enhanced data privacy and can seamlessly incorporate advanced, encrypted and secure computation techniques. Despite the increased computational requirements and reduced ability to explore and troubleshoot issues with the data (104), the benefits of FL may outweigh these additional costs.
Therefore, federated learning can potentially mitigate the risks associated with national regulations if implemented in a manner that is secure by design and by default. Its use may also make patients and institutions more confident in participating in clinical studies by reducing privacy and data misuse risks. However, the reliable use of federated learning and its effective translation into clinical practice require a concerted effort by machine learning, clinical, and information technology specialists. All their skills are necessary to accurately evaluate the associated risks and expand its practical applications in bioinformatics beyond the early-stage applications reviewed in this paper.
Conflict of interest statement
The authors declare that the research was conducted without any commercial or financial relationships that could potentially create a conflict of interest.
Author contributions
DM: Conceptualization, Funding acquisition, Investigation, Methodology, Writing - original draft, Writing - review & editing.
MS: Conceptualization, Investigation, Methodology, Writing - original draft, Writing - review & editing.
FG: Methodology, Writing - original draft, Writing - review & editing.
HS: Writing - original draft, Writing - review & editing.
AML: Funding acquisition, Project administration, Writing - original draft, Writing - review & editing.
IH: Writing - original draft, Writing - review & editing.
MF: Conceptualization, Funding acquisition, Project administration, Writing - original draft, Writing - review & editing.
JvS: Writing - original draft, Writing - review & editing.
SWvdL: Funding acquisition, Project administration, Writing - original draft, Writing - review & editing.